OpenStudy (anonymous):

Solve the following equation. Be sure to check the answer in the original equation if you multiply both sides by an expression that contains the variable. (Problem in comments...)

OpenStudy (anonymous):

\[\frac{ 15 }{ x^2-9 } +\frac{ 3 }{ x+3 } =\frac{ 2 }{ x-3 }\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Multiply all sides by \(x^2-9\).

OpenStudy (anonymous):

Notice that \(x^2-9=(x-3)(x+3)\).

OpenStudy (anonymous):

So we get \[ 15 + 3(x-3) = 2(x+3) \]

OpenStudy (anonymous):

That's the answer?

OpenStudy (anonymous):

you need to simplify and solve for \(x\), but the problem is much simpler now.

OpenStudy (anonymous):

is it, 15+3x+9=2x+6?

ganeshie8 (ganeshie8):

\(15 + 3(x-3) = 2(x+3) \) becomes \(15 + 3x \color{red}{-} 9= 2x+6 \) right ?

ganeshie8 (ganeshie8):

after that, you need to solve \(x\), which is same as isolating \(x\)

OpenStudy (anonymous):

Thats when i factor right?

OpenStudy (anonymous):

I am having some trouble with this... Did you work out the answer @cookiibabii93 Did you get x to be 0, because thats what i got...

OpenStudy (anonymous):

You should get \(x=0\).

OpenStudy (anonymous):

@wio I kind of get scared when x=0 because I always expect x to never be 0.